# Scaling Robust Optimization for Multi-Agent Robotic Systems: A Distributed Perspective

### v1

@@ Abstract
This paper presents a novel distributed robust optimization scheme for steering distributions of multi-agent systems under stochastic and deterministic uncertainty. Robust optimization is a subfield of optimization which aims in discovering an optimal solution that remains robustly feasible for all possible realizations of the problem parameters within a given uncertainty set. Such approaches would naturally constitute an ideal candidate for multi-robot control, where in addition to stochastic noise, there might be exogenous deterministic disturbances. Nevertheless, as these methods are usually associated with significantly high computational demands, their application to multi-agent robotics has remained limited. The scope of this work is to propose a scalable robust optimization framework that effectively addresses both types of uncertainties, while retaining computational efficiency and scalability. In this direction, we provide tractable approximations for robust constraints that are relevant in multi-robot settings. Subsequently, we demonstrate how computations can be distributed through an Alternating Direction Method of Multipliers (ADMM) approach towards achieving scalability and communication efficiency. Simulation results highlight the performance of the proposed algorithm in effectively handling both stochastic and deterministic uncertainty in multi-robot systems. The scalability of the method is also emphasized by showcasing tasks with up to 100 agents. The results of this work indicate the promise of blending robust optimization, distribution steering and distributed optimization towards achieving scalable, safe and robust multi-robot control.
@@ I Introduction
  • We provide simulation results that demonstrate the effectiveness of the proposed algorithm in handling stochastic and deterministic uncertainty. The scalability of the method is also highlighted by showing multi-robot tasks with up to 100 agents.
@@ II-C Dynamics
where the matrices ${\bf G}_{0}^{i},{\bf G}_{u}^{i},{\bf G}_{{\bm{w}}}^{i},{\bf G}_{\zeta}^{i}$ are provided in Section I of the Supplementary Material (SM).
@@ II-D Characterizations of Uncertainty
The unknown deterministic disturbances $\bm{\zeta}^{i}$ typically stem from external factors that cannot be accurately represented through stochastic signals. This creates a requirement to formulate optimization problems and achieve optimal policies that are feasible in a robust sense for all possible values of such disturbances in a bounded uncertainty set. Examples of such uncertainty sets include ellipsoids, polytopes, and others [5]. In this work, we assume the sequence $\bm{\zeta}^{i}$ to be lying inside an ellipsoidal uncertainty set defined as follows
@@ II-E Control Policies
$u^{i}_{k}=\bar{u}^{i}_{k}+\sum_{\ell=0}^{k}K^{i}_{k,\ell}\delta^{i}_{\ell},\;\forall\;k\in\llbracket 0,T-1\rrbracket,$ (9)
where $\bar{u}^{i}_{k}\in\mathbb{R}^{n_{u_{i}}}$ are the feed-forward control inputs and $K^{i}_{k,\ell}\in\mathbb{R}^{n_{u_{i}}\times n_{x_{i}}}$ are feedback gains on the purified states. The above policies can be written in compact form as
@@
$\bar{{\bm{u}}}^{i}=\begin{bmatrix}\bar{u}_{0}^{i};&\dots;&\bar{u}_{T-1}^{i}\end{bmatrix},$ (12)
${\bf K}^{i}=\begin{bmatrix}K_{0,0}^{i}&{\bf 0}&\dots&{\bf 0}&{\bf 0}\\ K_{1,0}^{i}&K_{1,1}^{i}&\dots&{\bf 0}&{\bf 0}\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ K_{T-1,0}^{i}&K_{T-1,1}^{i}&\dots&K_{T-1,T-1}^{i}&{\bf 0}\end{bmatrix}.$ (17)
@@ II-F Robust Constraints
where ${\bf A}_{i}\in\mathbb{R}^{m\times n_{x_{i}}}$, $\bm{b}_{i}\in\mathbb{R}^{m}$. In a robotics setting, such constraints could typically represent bounds on the mean position, velocity, etc.
@@
where $p$ is the probability threshold. It should be noted that these constraints are also affected by the stochastic noise since they probabilistic constraints on the entire state and not only on the state mean.
@@ III-A Reformulation of Semi-infinite Constraints
We start by analyzing the effect of the disturbances on the system dynamics. In particular, since the robust constraints (18) - (20) depend on $\mathbb{E}[{\bm{x}}^{i}]$, let us first examine the expression for $\mathbb{E}[{\bm{x}}^{i}]$ given by
@@
since $\mathbb{E}[{\bm{w}}^{i}]=0$ (derivation can be referenced from Section II of the SM). It is important to emphasize that the mean state relies only on the deterministic - and not the stochastic - disturbances. For convenience, we also define
@@ III-A1 Robust linear mean constraints
with $\bar{{\bm{a}}}_{i}\in\mathbb{R}^{n_{x_{i}}},\bar{b}_{i}\in\mathbb{R}$ and $\epsilon_{i}\in\mathbb{R}^{+}$, which would imply the following constraints
@@ III-A2 Robust nonconvex norm-of-mean constraints
Although the problem in the LHS is convex, the equivalent tractable constraints to the above, obtained by applying S-lemma [19], might not be convex due to the non-convexity of the overall constraint (details can be referenced from Section III of the SM). Thus, we first linearize the constraint (19) with respect to control parameters $\bar{{\bm{u}}}^{i},{\bf K}^{i}$ around the nominal values $\bar{{\bm{u}}}^{i,l},{\bf K}^{i,l}$. Through this approach, we obtain the following constraint
@@
where $\mathscr{Q}({\bf K}^{i},{\bf K}^{i,l}),\bar{\mathscr{Q}}({\bf K}^{i},{\bf K}^{i,l},\bar{{\bm{u}}}^{i},\bar{{\bm{u}}}^{i,l})$ and $q(\bar{{\bm{u}}}^{i},\bar{{\bm{u}}}^{i,l})$ are provided in Section III of the SM. The equivalent tractable constraints to the above constraint can be given as [19] -
@@
The derived equivalent tractable constraints can be used in cases where the number of robust non-convex constraints of the defined form is relatively small. However, in applications that involve a significant number of such constraints (e.g., collision avoidance constraints), this formulation might prove to be significantly expensive computationally. For instance, consider an obstacle avoidance scenario, where each agent $i$ needs to enforce the constraints of the above form for each time step. Then, each agent $i$ in an environment with $n_{obs}$ obstacles would have $n_{obs}T$ number of constraints of the form (35). Each constraint of the form (35) involves a matrix of size $\bar{n}_{i}+1$ (when $\bm{\Gamma}_{i}=I$, $\bar{n}_{i}=n_{x_{i}}+Tn_{d_{i}}$). Thus, the computational cost of the above formulation increases exponentially with the number of obstacles, time horizon, and dimension of the disturbances. In order to overcome this issue, we propose a second approach for deriving equivalent/approximate tractable versions of the constraint (19).
@@
In this approach, we consider the following tighter approximation of the above constraint
$\|{\bf A}_{i}{\bm{\mu}}_{x_{k},\bar{u}}^{i}-\bm{b}_{i}\|_{2}-\|{\bf A}_{i}\tilde{{\bf M}}_{k}^{i}{\bm{\zeta}}^{i}\|_{2}\geq c_{i},\quad\forall{\bm{\zeta}}^{i}\in\pazocal{U}_{i},$ (37)
@@
Let us now introduce a slack variable $\tilde{c}_{i}$ such that we split the above constraint into the following system of two constraints.
@@
Note however that the constraint (39) is non-convex. We address this by linearizing the constraint around the nominal trajectory $\bar{{\bm{u}}}^{i,l}$. In addition, the constraint (40) is still a semi-infinite constraint; thus, we need to derive its equivalent tractable constraints. Since the LHS in the constraint (40) involves a non-convex optimization problem, deriving the equivalent tractable constraints is NP-hard. Hence, we derive tighter approximations of the constraint (40) in the following proposition.
@@
The tighter approximate constraints to the semi-infinite constraint (40) are given by
@@
Provided in Section III of the SM. ∎
In this formulation, only the constraints (42) include ${\bf S}_{i}$. This formulation is useful in the case where there are multiple non-convex constraints of the form (19) at $x_{k}^{i}$, which have the same ${\bf A}_{i}$ matrix. In such a case, we only need to enforce the constraint (42) once per each unique ${\bf A}_{i}$ matrix in the constraints. For instance, let us again consider the collision avoidance example. Since the collision constraints (both obstacle and inter-agent) are applied to the agent’s position, the matrix ${\bf A}_{i}$ remains the same in all such constraints, irrespective of the number of obstacles. Therefore, it is only the amount of constraints of the form (41) that would increase with the number of obstacles.
@@
where $\tilde{{\bf A}}_{ij}=\begin{bmatrix}{\bf A}_{i},&-{\bf A}_{j}\end{bmatrix}$, $\tilde{{\bm{x}}}^{ij}_{k}=\begin{bmatrix}{\bm{x}}^{i}_{k};&{\bm{x}}^{j}_{k}\end{bmatrix}$. The uncertainty involved for the variable $\tilde{{\bm{x}}}^{ij}$ would be $\tilde{{\bm{\zeta}}}^{ij}=\begin{bmatrix}{\bm{\zeta}}^{i};&{\bm{\zeta}}^{j}\end{bmatrix}$. Further, it should be noted that if ${\bm{\zeta}}^{i},{\bm{\zeta}}^{j}$ belong to the defined ellipsoidal sets $\pazocal{U}_{i}$, $\pazocal{U}_{j}$ respectively, then the uncertainty vector $\tilde{{\bm{\zeta}}}^{ij}$ can also be modeled by an ellipsoidal set of the defined form (II-D). Thus, the constraint (43) has a similar form as (19), and applying the initial approach would result in semi-definite constraints that are of the size $\bar{n}_{i}+\bar{n}_{j}+1$. This might be particularly disadvantageous in a large-scale multi-agent setup, which might involve a significant number of neighbor agents. Thus, we derive equivalent/approximate tractable constraints using the second approach.
@@
$\|{\bf A}_{i}{\bm{\mu}}_{x_{k},\bar{u}}^{i}-{\bf A}_{j}{\bm{\mu}}_{x_{k},\bar{u}}^{j}+{\bf A}_{i}\tilde{{\bf M}}_{k}^{i}{\bm{\zeta}}^{i}-{\bf A}_{j}\tilde{{\bf M}}_{k}^{j}{\bm{\zeta}}^{j}\|_{2}\geq c_{ij},$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\forall{\bm{\zeta}}^{i}\in\pazocal{U}_{i},~{}{\bm{\zeta}}^{j}\in\pazocal{U}_{j}.$ (44)
We then consider the following tighter approximation to the above constraint
$\|{\bf A}_{i}{\bm{\mu}}_{x_{k},\bar{u}}^{i}-{\bf A}_{j}{\bm{\mu}}_{x_{k},\bar{u}}^{j}\|_{2}-\|{\bf A}_{i}\tilde{{\bf M}}_{k}^{i}{\bm{\zeta}}^{i}-{\bf A}_{j}\tilde{{\bf M}}_{k}^{j}{\bm{\zeta}}^{j}\|_{2}\geq c_{ij}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\forall{\bm{\zeta}}^{i}\in\pazocal{U}_{i},~{}{\bm{\zeta}}^{j}\in\pazocal{U}_{j}.$ (45)
and similarly with the previous case, we introduce a slack variable $\tilde{c}_{ij}$, through which we can rewrite the above constraint as follows
$\|{\bf A}_{i}{\bm{\mu}}_{x_{k},\bar{u}}^{i}-{\bf A}_{j}{\bm{\mu}}_{x_{k},\bar{u}}^{j}\|_{2}\geq\tilde{c}_{ij},$ (46)
@@
The tighter approximate constraints to the semi-infinite constraint (47) are given by
@@
It should be noted that the constraints (49), (50) will be the same as (42), if the ${\bf A}_{i}$ matrix is the same (e.g., collision avoidance scenarios), therefore alleviating the need to introduce them as additional constraints in our framework.

### v2

@@ Abstract
This paper presents a novel distributed robust optimization scheme for steering distributions of multi-agent systems under stochastic and deterministic uncertainty. Robust optimization is a subfield of optimization which aims to discover an optimal solution that remains robustly feasible for all possible realizations of the problem parameters within a given uncertainty set. Such approaches would naturally constitute an ideal candidate for multi-robot control, where in addition to stochastic noise, there might be exogenous deterministic disturbances. Nevertheless, as these methods are usually associated with significantly high computational demands, their application to multi-agent robotics has remained limited. The scope of this work is to propose a scalable robust optimization framework that effectively addresses both types of uncertainties, while retaining computational efficiency and scalability. In this direction, we provide tractable approximations for robust constraints that are relevant in multi-robot settings. Subsequently, we demonstrate how computations can be distributed through an Alternating Direction Method of Multipliers (ADMM) approach towards achieving scalability and communication efficiency. All improvements are also theoretically justified by establishing and comparing the resulting computational complexities. Simulation results highlight the performance of the proposed algorithm in effectively handling both stochastic and deterministic uncertainty in multi-robot systems. The scalability of the method is also emphasized by showcasing tasks with up to hundreds of agents. The results of this work indicate the promise of blending robust optimization, distribution steering and distributed optimization towards achieving scalable, safe and robust multi-robot control.
@@ I Introduction
  • All computational improvements are theoretically justified by comparing the resulting computational complexities of different potential formulations.
  • Simulation results that demonstrate the effectiveness of the proposed algorithm in handling stochastic and deterministic uncertainty are provided in a variety of multi-robot tasks. The scalability of the method is also highlighted by showcasing scenarios with up to hundreds of agents and by comparing against a centralized approach.
@@ II-C Dynamics
where the matrices ${\bf G}_{0}^{i},{\bf G}_{u}^{i},{\bf G}_{{\bm{w}}}^{i},{\bf G}_{\zeta}^{i}$ are provided in the Appendix-A.
@@ II-D Characterizations of Uncertainty
The unknown deterministic disturbances $\bm{\zeta}^{i}$ typically stem from external factors that cannot be accurately represented through stochastic signals. This creates a requirement to formulate optimization problems and achieve optimal policies that are feasible in a robust sense for all possible values of such disturbances in a bounded uncertainty set. In this work, we assume the sequence $\bm{\zeta}^{i}$ to be lying inside an ellipsoidal uncertainty set defined as follows
@@ II-E Control Policies
$u^{i}_{k}=\bar{u}^{i}_{k}+\sum_{\ell=k-\gamma_{h}+1}^{k}K^{i}_{k,\ell}\delta^{i}_{\ell},\;\forall\;k\in\llbracket 0,T-1\rrbracket,$ (9)
where $\bar{u}^{i}_{k}\in\mathbb{R}^{n_{u_{i}}}$ are the feed-forward control inputs, and $K^{i}_{k,\ell}\in\mathbb{R}^{n_{u_{i}}\times n_{x_{i}}}$ are feedback gains on the purified states. Further, $\gamma_{h}$ defines the length of the history interval considered for the feedback. The above policies can be written in compact form (when $\gamma_{h}=T$) as
@@
$\bar{{\bm{u}}}^{i}=\begin{bmatrix}\bar{u}_{0}^{i};&\dots;&\bar{u}_{T-1}^{i}\end{bmatrix},$
${\bf K}^{i}=\begin{bmatrix}K_{0,0}^{i}&{\bf 0}&\dots&{\bf 0}&{\bf 0}\\ K_{1,0}^{i}&K_{1,1}^{i}&\dots&{\bf 0}&{\bf 0}\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ K_{T-1,0}^{i}&K_{T-1,1}^{i}&\dots&K_{T-1,T-1}^{i}&{\bf 0}\end{bmatrix}.$
@@ II-F Robust Constraints
where ${\bf A}_{i}\in\mathbb{R}^{m\times(T+1)}n_{x_{i}}$, $\bm{b}_{i}\in\mathbb{R}^{m}$. In a robotics setting, such constraints could typically represent bounds on the mean position, velocity, etc.
@@
where ${\bm{a}}_{i}\in\mathbb{R}^{(T+1)n_{x_{i}}}$, $b_{i}\in\mathbb{R}$, $p$ is the probability threshold. It should be noted that these constraints are also affected by the stochastic noise since those are probabilistic constraints on the entire state and not only on the state mean.
@@ III-A Reformulation of Semi-infinite Constraints
We start by analyzing the effect of the disturbances on the system dynamics. In particular, since the robust constraints (11) 13) (13) depend on $\mathbb{E}[{\bm{x}}^{i}]$, let us first examine the expression for $\mathbb{E}[{\bm{x}}^{i}]$ given by
@@
since $\mathbb{E}[{\bm{w}}^{i}]=0$ (see Appendix-B). It is important to emphasize that the mean state relies only on the deterministic - and not the stochastic - disturbances. For convenience, let us also define
@@ III-A1 Robust linear mean constraints
with $\bar{{\bm{a}}}_{i}\in\mathbb{R}^{(T+1)}n_{x_{i}},\bar{b}_{i}\in\mathbb{R}$ and $\epsilon_{i}\in\mathbb{R}^{+}$, which would imply the following constraints
@@ III-A2 Robust nonconvex norm-of-mean constraints
Although the problem in the LHS is convex, the equivalent tractable constraints to the above, obtained by applying S-lemma [20], might not be convex due to the non-convexity of the overall constraint (details can be referenced from Appendix-E). Thus, we first linearize the constraint (12) with respect to control parameters $\bar{{\bm{u}}}^{i},{\bf K}^{i}$ around the nominal values $\bar{{\bm{u}}}^{i,l},{\bf K}^{i,l}$. Through this approach, we obtain the following constraint
@@
where $\mathscr{Q}({\bf K}^{i},{\bf K}^{i,l}),\bar{\mathscr{Q}}({\bf K}^{i},{\bf K}^{i,l},\bar{{\bm{u}}}^{i},\bar{{\bm{u}}}^{i,l})$ and $q(\bar{{\bm{u}}}^{i},\bar{{\bm{u}}}^{i,l})$ are provided in Appendix-E. The equivalent tractable constraints to the above constraint can be given as [20] -
@@
The derived equivalent tractable constraints can be used in cases where the number of robust non-convex constraints of the defined form is relatively small. However, in applications that involve a significant number of such constraints (e.g., collision avoidance constraints), this formulation might prove to be significantly expensive computationally. For instance, consider an obstacle avoidance scenario, where each agent $i$ needs to enforce the constraints of the above form for each time step. Then, each agent $i$ in an environment with $n_{obs}$ obstacles would have $n_{obs}T$ number of constraints of the form (28). Each constraint of the form (28) involves a matrix of size $\bar{n}_{i}+1$ (when $\bm{\Gamma}_{i}=I$, $\bar{n}_{i}=n_{x_{i}}+Tn_{d_{i}}$ ). Thus, the computational cost of the above formulation increases exponentially with the number of obstacles, time horizon, and the dimension of the disturbances (see Section III-D for a rigorous complexity analysis). In order to overcome this issue, we propose a second approach for deriving equivalent/approximate tractable versions of the constraint (12).
@@
Based on the triangle inequality, we consider the following tighter approximation of the above constraint
$\|{\bf A}_{i}{\bm{\mu}}_{x_{k},\bar{u}}^{i}-\bm{b}_{i}\|_{2}-\|{\bf A}_{i}\tilde{{\bf M}}_{k}^{i}{\bm{\zeta}}^{i}\|_{2}\geq c_{i},\quad\forall{\bm{\zeta}}^{i}\in\pazocal{U}_{i}$ (30)
@@
Let us now introduce a slack variable $\tilde{c}_{i}$ such that we split the above constraint into the following system of two constraints
@@
Note, however, that the constraint (32) is non-convex. We address this by linearizing the constraint around the nominal trajectory $\bar{{\bm{u}}}^{i,l}$. In addition, the constraint (33) is still a semi-infinite constraint; thus, we need to derive its equivalent tractable constraints. Since the LHS in the constraint (33) involves a non-convex optimization problem, deriving the equivalent tractable constraints is NP-hard. Hence, we derive tighter approximations of the constraint (33) in the following proposition.
@@
The tighter approximate constraints to the semi-infinite constraint (33) are given by the following second-order cone constraints (SOCP constraints)
@@
Provided in Appendix-F. ∎
In this formulation, only the constraints (35) include ${\bf S}_{i}$. This formulation is useful in the case where there are multiple non-convex constraints of the form (12) at $x_{k}^{i}$, which have the same ${\bf A}_{i}$ matrix. In such a case, we only need to enforce the constraint (35) once per each unique ${\bf A}_{i}$ matrix in the constraints. For instance, let us again consider the collision avoidance example. Since the collision constraints (both obstacle and inter-agent) are applied to the agent’s position, the matrix ${\bf A}_{i}$ remains the same in all such constraints, irrespective of the number of obstacles. Therefore, it is only the amount of constraints of the form (34) that would increase with the number of obstacles. Further, the geometrical interpretation of the tighter approximations proposed for a 2D case can be found in Appendix-H.
@@
where $\tilde{{\bf A}}_{ij}=\begin{bmatrix}{\bf A}_{i},&-{\bf A}_{j}\end{bmatrix}$, $\tilde{{\bm{x}}}^{ij}_{k}=\begin{bmatrix}{\bm{x}}^{i}_{k};&{\bm{x}}^{j}_{k}\end{bmatrix}$. The uncertainty involved for the variable $\tilde{{\bm{x}}}^{ij}$ would be $\tilde{{\bm{\zeta}}}^{ij}=\begin{bmatrix}{\bm{\zeta}}^{i};&{\bm{\zeta}}^{j}\end{bmatrix}$. Further, it should be noted that if ${\bm{\zeta}}^{i},{\bm{\zeta}}^{j}$ belong to the defined ellipsoidal sets $\pazocal{U}_{i}$, $\pazocal{U}_{j}$ respectively, then the uncertainty vector $\tilde{{\bm{\zeta}}}^{ij}$ can also be modeled by an ellitope (which is an intersection of ellipsoids). Thus, by using the initial approach, the constraint (36) would result in semi-definite constraints that are of the size $\bar{n}_{i}+\bar{n}_{j}+1$ [20]. This might be particularly disadvantageous in a large-scale multi-agent setup, which might involve a significant number of neighbor agents. Thus, we derive equivalent/approximate tractable constraints using the second approach.
@@
$\forall{\bm{\zeta}}^{i}$ $\in\pazocal{U}_{i},~{}{\bm{\zeta}}^{j}\in\pazocal{U}_{j},$
$\|{\bf A}_{i}{\bm{\mu}}_{x_{k},\bar{u}}^{i}-{\bf A}_{j}{\bm{\mu}}_{x_{k},\bar{u}}^{j}+{\bf A}_{i}\tilde{{\bf M}}_{k}^{i}{\bm{\zeta}}^{i}-$ ${\bf A}_{j}\tilde{{\bf M}}_{k}^{j}{\bm{\zeta}}^{j}\|_{2}\geq c_{ij}.$
We then consider the following tighter approximation to the above constraint-
$\forall{\bm{\zeta}}^{i}$ $\in\pazocal{U}_{i},~{}{\bm{\zeta}}^{j}\in\pazocal{U}_{j},$
[…]
and similarly, with the previous case, we introduce a slack variable $\tilde{c}_{ij}$, through which we can rewrite the above constraint as follows
$\|{\bf A}_{i}{\bm{\mu}}_{x_{k},\bar{u}}^{i}-{\bf A}_{j}{\bm{\mu}}_{x_{k},\bar{u}}^{j}\|_{2}\geq\tilde{c}_{ij}$ (37)
[…]
@@
The tighter approximate constraints to the semi-infinite constraint (38) are given by the following second-order cone constraints (SOCP)
@@
Proof.
Provided in Appendix-G. ∎
It should be noted that the constraints (40), (41) will be the same as (35) if the ${\bf A}_{i}$ matrix is the same (e.g., collision avoidance scenarios), therefore alleviating the need to introduce them as additional constraints in our framework.